OpenStudy (anonymous):

how does x^2-4 meet the requirements of a function with two real rational solutions?

OpenStudy (solomonzelman):

If you set it equal zero, and solve for x you will see. \[x^2-4=0\]\[x^2-4+4=0+4\]\[x^2=4\]\[x=±\sqrt{4}\]\[x=±2\]to write the funtion out \[f(x)=(x+2)(x-2)\]\[f(x)=x^2-4\]